# Neural Networks can detect model-free static arbitrage strategies

### v1

@@ Abstract.
July 13, 2023
${}^{1}$ Singapore, Division of Mathematical Sciences,
@@
${}^{2}$ational University of Singapore, Department of Mathematics,
@@ 2.1. Setting
Let $(K,\pi)\in[0,\overline{K}]^{N}\times[0,\overline{\pi}]^{2N}$. Then, we call a static trading strategy $(a,h)\in[\underline{a},\infty)\times[0,\overline{H}]^{2N}$ a model-free static arbitrage strategy if the following two conditions hold.
@@ 2.2. Main results
The main idea to derive Theorem 2.5 and Theorem 2.6 relies on the relation between arbitrage and supehedging of the 00-payoff function. The following result establishes that for any prescribed $\varepsilon>0$ there exists a single neural network such that for any given strikes $K=(K_{i,j})_{i=1,\dots,N_{\Psi},\atop j=1,\dots,n_{i}}$ and option prices $\pi=(\pi_{i,j}^{+},\pi_{i,j}^{-})_{i=1,\dots,N_{\Psi},\atop j=1,\dots,n_{i}}$ defining the market, the neural network produces a static trading strategy which superhedges the 00-payoff for all possible values $S\in\mathcal{S}$ whose price is $\varepsilon$-optimal.
@@ 3.1.1. Training with data of the S&P 500
Figure 1. The histogram shows the distribution of the net profit $\mathcal{I}_{S_{i,j}}(K_{i},a_{i},h_{i})-f(\pi_{i},a_{i},h_{i})$ for $i=1,\dots,5000$, $j=1,\dots,200$ of the strategy trained as described in Section 3.1.1.
@@ 3.1.2. Backtesting with historical option prices
We apply the strategy trained in Section 3.1.1 to the prices observed on each of the $33$ trading days and evaluate it on the realized values of the $5$ underlying securities at maturity. In Table 2 and Figure 2 we summarize the net profits of the $33$ strategies. Note that to apply the trained neural network from Section 3.1.1, we first scale all the financial instruments such that the spot values of the underlying securities equal 1, as described in Section 3.1.1. Then, after applying the strategies to the scaled inputs, we rescale the values of the involved quantities back to unnormalized values, and we report in Table 2 and Figure 2 the net profits for both cases: after rescaling the values of the underlying securities, options, and strikes to unnormalized values, as well as without scaling back. The results of the backtesting study reveal that even though the neural network from Section 3.1.1 was trained on data extracted at a different day ($25$ April $2025$) involving call options with a different maturity written on other assets, the resultant strategy still allows to trade profitably in the majority of cases, showcasing the robustness of our algorithm.
@@ 4. Approximation of optimal solutions of general convex semi-infinite programs by neural networks
In this section we show for a certain class of convex semi-infinite optimization problems (CSIP) that each of them can be approximately solved by a single neural networks. More precisely, for every prescribed accuracy $\varepsilon>0$ we show that there exists a single neural network which outputs a feasible solution which is $\varepsilon$-optimal. This class of convex semi-infinite problems covers the setting of static arbitrage detection introduced in Section 2 as special case. We leave further applications for future research.
@@ 4.1. Setting
    ($\mathcal{I}{}_{s}(x,\overline{a}^{\operatorname{UB}},y)$ $=\mathcal{I}_{s}(x,\overline{a}^{\operatorname{UB}},y)-\mathcal{I}_{s}(x,\underline{a},y)+\mathcal{I}_{s}(x,\underline{a},y)$
@@ 5.2.1. Auxiliary Results
Now we consider the case that $a>\overline{a}^{\operatorname{UB}}$. Note that in this case $\mathcal{I}_{s}(x,a,y)>0$ for all $s\in\mathcal{S}$ due to the strict monotonocity of $\mathcal{I}_{s}$ and by Remark 4.4 (i). Hence, by the continuity of $\mathcal{I}_{s}$, there exists some $n_{0}\in\mathbb{N}$ such that for all $n\geq n_{0}$ we have $\mathcal{I}_{s}(x^{(n)},a,y)>0$ implying that $(a,y)\in\Gamma_{\overline{a}}(x^{(n)})$ for all $n\geq n_{0}$. Thus, we conclude with (Aliprantis, Theorem 17.21) the lower hemicontinuity of the map from (5.5) also for the case $a>\overline{a}^{\operatorname{UB}}$.

### v2

@@ Abstract.
August 13, 2024
1NTU Singapore, Division of Mathematical Sciences,
@@
2National University of Singapore, Department of Mathematics,
@@ 2.1. Setting
Let $(K,\pi)\in[0,\overline{K}]^{N}\times[0,\overline{\pi}]^{2N}$. We call a static trading strategy $(a,h)\in[\underline{a},\infty)\times[0,\overline{H}]^{2N}$ a model-free static arbitrage strategy if the following two conditions hold.
@@ 2.2. Main results
The main idea to derive Theorem 2.5 and Theorem 2.6 relies on the relation between arbitrage and superhedging of the 00-payoff function. The following result establishes that for any prescribed $\varepsilon>0$ there exists a single neural network such that for any given strikes $K=(K_{i,j})_{i=1,\dots,N_{\Psi},\atop j=1,\dots,n_{i}}$ and option prices $\pi=(\pi_{i,j}^{+},\pi_{i,j}^{-})_{i=1,\dots,N_{\Psi},\atop j=1,\dots,n_{i}}$ defining the market, the neural network produces a static trading strategy which superhedges the 00-payoff for all possible values $S\in\mathcal{S}$ whose price is $\varepsilon$-optimal.
@@ 3. The Numerics of Static Arbitrage Detection in Financial Markets
Remark 3.1.
Algorithm 1 is designed to fulfill two tasks simultaneously: first, to detect whether there exists arbitrage in the market (through (3.1)) and second how to exploit arbitrage if existent, while the focus through the design of the objective function lies on maximizing the profit if arbitrage exists. For the pure classification task of deciding whether arbitrage exists (without learning the associated arbitrage strategy), there exist better suited classification algorithms such as Logistic Regression (kleinbaum2002logistic), random forests (ho1995random) or XGBoost (chen2016xgboost). We leave the exploration of optimizing this related but still different task for future research.
@@ 3.1.1. Training with data of the S&P 500
[…]
Figure 1. The loss function as well as the training and test set accuracy in dependenceof the number of trained epochs.
[…]
All samples Conditional on wrong identification
count 1 000 000 108 000
mean 0.477996 0.468450
std 0.322428 0.323814
min -0.179760 -0.101957
25% 0.215138 0.202838
50% 0.411158 0.398024
75% 0.705576 0.700997
max 3.086741 2.980903
[…]
Figure 2. Left: The histogram shows the distribution of the net profit $\mathcal{I}_{S_{i,j}}(K_{i},a_{i},h_{i})-f(\pi_{i},a_{i},h_{i})$ for $i=1,\dots,5000$, $j=1,\dots,200$ of the strategy trained as described in Section 3.1.1. Right: The histogram shows the distribution of the net profit conditional on wrong identification of arbitrage.
3.1.2. Stability of the results and choice of hyperparameters
[…]
@@ 3.1.2. Stability of the results and choice of hyperparameters
Figure 3. The histogram shows the distribution of the net profit $\mathcal{I}_{S_{i,j}}(K_{i},a_{i},h_{i})-f(\pi_{i},a_{i},h_{i})$ for $i=1,\dots,5000$, $j=1,\dots,200$ of the strategy trained as described in Section 3.1.1 but on a reduced and balanced training set where 50 % of the sampled constitute arbitrage situations.
Next, we study the consequences of varying the used hyperparameters and we report in Table 3 the share of correct predictions, precision, recall, and F1 score for different configurations evaluated with respect to arbitrage detection on the test set. Moreover, Table 4 shows the performance of the profit of the trained strategies on the test set in dependence of different choices of hyperparameters. Taken together, the results reveal that the results do not vary significantly across different hyperparameter-configurations while a learning rate of $0.0001$, a depth of $5$ for the neural network, and no regularization seem to be close-to-optimal hyperparameter choices. Note also that the choice of a smaller learning rate of $0.001$ leads to a larger mean profit which however comes at the cost of a much higher standard deviation indicating a suboptimal solution. Moreover, increasing the depth of the neural network to $10$ prevents the neural network from learning a similar profitable strategy.
Learning Rate, Depth, Regularization Correct Predictions Precision Recall $F_{1}$ Score
0.0001,2,none 0.8216 0.884871 0.878796 0.881823
0.0001,3,none 0.8228 0.866195 0.905994 0.885648
0.0001,4,none 0.8606 0.920065 0.893583 0.906631
0.0001,5,none 0.8920 0.955911 0.898865 0.926511
0.0001,10,none 0.2510 0.937500 0.011883 0.023468
0.001,5,none 0.7186 0.811518 0.818590 0.815039
1e-05,5,none 0.6936 0.866667 0.703723 0.776741
0.0001,5,l1 0.8936 0.959097 0.897808 0.927441
0.0001,5,l2 0.8186 0.815928 0.982044 0.891312
Table 3. Performance metrics for different model parameters, evaluated on the test set.
Parameters Count Mean Std Min 25% 50% 75% Max
0.0001,2,none 1000000 0.439912 0.297357 -0.267825 0.203189 0.367542 0.644060 3.132133
0.0001,3,none 1000000 0.436585 0.302734 -0.209222 0.193420 0.360132 0.648741 3.113397
0.0001,4,none 1000000 0.466350 0.310355 -0.132638 0.217923 0.397229 0.681759 3.597595
0.0001,5,none 1000000 0.477481 0.321754 -0.157909 0.215212 0.410469 0.704891 3.220638
0.0001,10,none 1000000 0.034884 0.025370 -0.008109 0.015160 0.031994 0.050856 0.182104
0.001,5,none 1000000 1.363648 0.837385 -0.716115 0.733738 1.260230 1.896849 5.774780
1e-05,5,none 1000000 0.406486 0.181477 -0.189432 0.274503 0.392536 0.526439 1.299631
0.0001,5,l1 1000000 0.464646 0.317831 -0.158561 0.204477 0.395615 0.690994 3.072829
0.0001,5,l2 1000000 0.468799 0.320785 -0.174531 0.206054 0.399411 0.698279 3.331130
Table 4. Summary statistics for the net profit for strategies being trained using different hyperparameter combinations consisting of learning rate, depth of the neural network, and the type of regularization (None, $L_{1}$ or $L_{2}$), evaluated on the test set.
It is also noteworthy that the results are relatively robust with respect to the size of the training set on which the agent is trained. Indeed, in Table 5 we report the profits on the test set for agents that have been trained on training sets of varying size. The results show that for all considered sample sizes between $5000$ and $45000$ the resultant profit on the test set is on a similar scale.
Training set size 5000 10000 15000 20000 25000 30000 35000 40000 45000
Count 1000000 1000000 1000000 1000000 1000000 1000000 1000000 1000000 1000000
Mean 0.546421 0.449859 0.459965 0.458827 0.466374 0.568733 0.538420 0.468137 0.477996
Std 0.327194 0.304473 0.315945 0.309333 0.313317 0.247745 0.238015 0.321906 0.322428
Min -0.175621 -0.177209 -0.214879 -0.177739 -0.188140 -0.133989 -0.112773 -0.215613 -0.179760
25% 0.292485 0.198430 0.201478 0.205287 0.210050 0.407968 0.386002 0.209109 0.215138
50% 0.493732 0.381197 0.389149 0.391185 0.398492 0.545277 0.514971 0.396287 0.411158
75% 0.763128 0.672535 0.686655 0.681570 0.691300 0.694669 0.655317 0.691226 0.705576
Max 3.851340 2.628088 3.276558 3.169339 3.166770 4.146131 3.460321 4.275422 3.086741
Table 5. Summary statistics for different sample sizes
Moreover, even though the optimization procedure is random due to the use of random initialization of weights and the use of the Adam optimizer, the resultant profits on the test turn out to be on the same scale when being trained several times with the same hyperparameter configuration, compare also Table 6 where we show the net profit for the optimization procedure as specified in Section 3.1.1 trained $10$ times independently.
Count 1000000 1000000 1000000 1000000 1000000 1000000 1000000 1000000 1000000 1000000
Mean 0.452637 0.472013 0.442239 0.474521 0.485375 0.449283 0.437489 0.464786 0.467977 0.461433
Std 0.309939 0.323847 0.306825 0.312429 0.327781 0.305616 0.311088 0.322367 0.309843 0.312393
Min -0.215139 -0.163364 -0.163844 -0.241484 -0.172754 -0.196785 -0.128142 -0.127955 -0.162520 -0.159971
25% 0.199371 0.211284 0.190744 0.227038 0.227661 0.199840 0.188147 0.210150 0.217428 0.205545
50% 0.380630 0.399340 0.369056 0.405421 0.413370 0.378361 0.354790 0.389295 0.399316 0.392809
75% 0.675014 0.697409 0.663718 0.688814 0.705762 0.668901 0.653276 0.680747 0.687475 0.685806
Max 3.252239 3.669023 2.746479 5.333844 4.395148 2.827598 3.379474 4.312405 3.028918 2.850765
Table 6. Summary statistics of the net profit on the test set for several strategies trained with the same hyperparameters.
@@ 3.1.3. Backtesting with historical option prices
We apply the strategy trained in Section 3.1.1 to the prices observed on each of the $33$ trading days and evaluate it on the realized values of the $5$ underlying securities at maturity. In Table 7 and Figure 4 we summarize the net profits of the $33$ strategies. Note that to apply the trained neural network from Section 3.1.1, we first scale all the financial instruments such that the spot values of the underlying securities equal 1, as described in Section 3.1.1. Then, after applying the strategies to the scaled inputs, we rescale the values of the involved quantities back to unnormalized values, and we report in Table 7 and Figure 4 the net profits for both cases: after rescaling the values of the underlying securities, options, and strikes to unnormalized values, as well as without scaling back. The results of the backtesting study reveal that even though the neural network from Section 3.1.1 was trained on data extracted at a different day ($25$ April $2023$) involving call options with a different maturity written on other assets, the resultant strategy still allows to trade profitably in the majority of cases, showcasing the robustness of our algorithm.
@@ 4. Approximation of optimal solutions of general convex semi-infinite programs by neural networks
In this section we show for a certain class of convex semi-infinite optimization problems (CSIP) that each of them can be approximately solved by a single neural network. More precisely, for every prescribed accuracy $\varepsilon>0$ we show that there exists a single neural network which outputs a feasible solution which is $\varepsilon$-optimal. This class of convex semi-infinite problems covers the setting of static arbitrage detection introduced in Section 2 as special case. We leave further applications for future research.
@@ 4.1. Setting
    $\mathcal{I}_{s}(x,\overline{a}^{\operatorname{UB}},y)$ $=\mathcal{I}_{s}(x,\overline{a}^{\operatorname{UB}},y)-\mathcal{I}_{s}(x,\underline{a},y)+\mathcal{I}_{s}(x,\underline{a},y)$
@@ 5.2.1. Auxiliary Results
Now we consider the case that $a>\overline{a}^{\operatorname{UB}}$. Note that in this case $\mathcal{I}_{s}(x,a,y)>0$ for all $s\in\mathcal{S}$ due to the strict monotonicity of $\mathcal{I}_{s}$ and by Remark 4.4 (i). Hence, by the continuity of $\mathcal{I}_{s}$, there exists some $n_{0}\in\mathbb{N}$ such that for all $n\geq n_{0}$ we have $\mathcal{I}_{s}(x^{(n)},a,y)>0$ implying that $(a,y)\in\Gamma_{\overline{a}}(x^{(n)})$ for all $n\geq n_{0}$. Thus, we conclude with (Aliprantis, Theorem 17.21) the lower hemicontinuity of the map from (5.5) also for the case $a>\overline{a}^{\operatorname{UB}}$.